OpenStudy (anonymous):

(4x^2 + 5x)(12x^2-5x-3)=0 for x

OpenStudy (anonymous):

\[(4x^2 + 5x)(12x^2-5x-3)=0\] need some factoring

OpenStudy (anonymous):

\[x (3 x+1) (4 x-3) (4 x+5) = 0\] now it should be easier

OpenStudy (anonymous):

thankee-sai. now it makes sense.

OpenStudy (anonymous):

x=0 x=-1/3 x=-5/4 x=3/4